The half life period of a first order reaction is 100 minute.calculate the time required for the reaction to be completed by 90%

Grade:12

1 Answers

Arun
6 years ago
Dear student
K = 0.693/t(1/2)
K = 0.693/100 = 0.00693
 
K = (2.303/t) log (a/a-x)
0.00693 = (2.303/t) log (100/10)
t = 2.303/.00693
t = 332.32 minutes
Regards
Arun (askIITians forum expert)